OpenStudy (anonymous):

help pls....Laplace transform: L{cos^3 (at)}

OpenStudy (anonymous):

Reduce the power of the trig function first: \[\cos^3at=\cos at\cos^2at\] Half-angle identity: \[\begin{align*}\cos^3at&=\cos at\left(\frac{1}{2}+\frac{1}{2}\cos2at\right)\\ &=\frac{1}{2}\cos at+\frac{1}{2}\cos at\cos 2at \end{align*}\] Recall the angle sum/difference identities: \[\begin{align*}\cos(x+y)&=\cos x\cos y-\sin x\sin y\\ \cos(x-y)&=\cos x\cos y+\sin x\sin y\end{align*}\] Adding the above equations, you have \[\begin{align*}\cos(x+y)+\cos(x-y)&=2\cos x\cos y\\ \\\frac{\cos(x+y)+\cos(x-y)}{2}&=\cos x\cos y \end{align*}\] So, you have \[\begin{align*}\cos^3at&=\frac{1}{2}\cos at+\frac{1}{2}\cos at\cos 2at\\\\ &=\frac{1}{2}\cos at+\frac{1}{2}\left(\frac{\cos3at+\cos at}{2}\right)\\\\ &=\frac{1}{2}\cos at+\frac{1}{4}\cos3at+\frac{1}{4}\cos at\\\\ &=\frac{3}{4}\cos at+\frac{1}{4}\cos3at\end{align*}\]

OpenStudy (anonymous):

Taking the LP should be easy if you have the formula for cosine.